Suppose that the proportions of blood phenotypes in a particular population are as follows A B AB o 0.46 0.07 0.04 0.43 Assuming that the phenotypes of two randomly selected Individuals are independent of one another, what is the probability that both phenotypes are O? (Enter your answer to four decimal places.) What is the probability that the phenotypes of two randomly selected individuals match? (Enter your answer to four decimal places.) Need Help?
Answer #1

The probability that both phenotypes are O type is obtained as:

P(both O) = P(0) * P(O) = 0.43 * 0.43 = 0.1849

There is a 18.49 percent chance that both the randomly selected individuals are of O phenotypes.

The probability that the phenotypes of two randomly selected individuals match is

P(Both match) P(A)*P(A)+P(B)*P(B)+P(AB)* P(AB)+P(O)*P(O) 0.46 0.46+0.07 0.070.040.04+0.43 0.43 0.4030

There is a 40.30 percent chance that both the randomly selected individuals match.